OpenStudy (anonymous):

A piece of metal with a mass of 130. g is placed in a 50 mL graduated cylinder. The water level rises from 20. mL to 41 mL. What is the density of the metal?

OpenStudy (saifoo.khan):

Density = mass/volume.

OpenStudy (saifoo.khan):

Density = 130/(41-20) Density = 130/ 21 = 6.19g/ml^3